# Orthogonalized Policy Optimization Decoupling Sampling Geometry from Optimization Geometry in RLHF

### v2

@@ Abstract
We propose a simple but structural remedy: formulate alignment as an orthogonal mirror descent problem, in which sampling geometry enters only as a linear driving force, while optimization geometry is determined independently by a mirror map. This perspective leads to a new alignment objective—Orthogonalized Policy Optimization (OPO)—obtained by choosing a Euclidean mirror map in likelihood-ratio space. The resulting objective admits a closed-form solution, linear (non-saturating) gradient dynamics, and a well-conditioned trust region, while remaining fully compatible with standard LLM training pipelines.
@@ 1 Introduction
  • We propose the Orthogonal Mirror Descent framework, which theoretically decouples these design axes.
@@ 2 Related Work
3 Theoretical Framework: Two Independent Design Axes
We formalize the OPO framework by decomposing the alignment objective into two orthogonal components.
3.1 Coordinates: Ratio Space vs. Log-Ratio Space
@@ 4.1 The General Framework
4.2 Axis 1: Sampling Geometry Specification
The sampling distribution $\rho_{\alpha}$ controls the exploit-explore tradeoff in the data. We define the preference weight $\omega_{\alpha}(y)$ generally as an $\alpha$-parameterized linear signal that reweights the advantage $A(y)$.
$\omega_{\alpha}(y)\propto\text{Projection}_{\alpha}(A(y),\pi_{\mathrm{ref}}(y))$ (4)
Specific constructions include the $\alpha$-divergence optimal target or other heavy-tailed reweighting schemes. This explicit definition avoids ambiguity: $\omega_{\alpha}$ represents the abstract "force" pulling the policy towards high-reward regions, decoupled from the optimization dynamics.
@@ 4.3 Axis 2: Optimization Geometry: Euclidean Mirror Map
This choice transforms the alignment problem into a quadratic optimization in ratio space.
4.4 Orthogonalized Policy Optimization (OPO)
Substituting the Euclidean geometry into our framework yields the OPO objective:
@@ 4.4 Orthogonalized Policy Optimization (OPO)
where $\mu=1/\eta$ is the stiffness parameter. The solution is available in closed form: $v^{*}(y)=\omega_{\alpha}(y)/\mu$.
@@
A physical analogy elucidates the advantages of the Euclidean geometry.
  • Ratio Space as Amplitude: In quantum mechanics, superpositions are linear in amplitude ($\psi$). Similarly, OPO operates in the ratio space $v$, allowing preference signals to superpose linearly without saturation.
  • KL Space as Probability: KL-based methods operate in log-probability space. Analogous to adding energies in a Boltzmann distribution, this enforces exponential scaling, which leads to "probability collapse" or saturation when signals are strong.
By working in "amplitude" (ratio) space with quadratic "energy" (regularization), OPO maintains a non-saturating linear response field, crucial for long-chain reasoning where we must avoid premature collapse to a single path.
4.5 Log-Ratio Approximation
For implementation in LLMs (which output logits), we use the log-ratio $\Delta_{\theta}=\log\pi_{\theta}-\log\pi_{\mathrm{ref}}$. As shown in Appendix A, for small trust regions, $v\approx\Delta$.
$\boxed{\mathcal{L}_{\text{OPO}}^{\text{log}}\approx-\sum_{y\in S}\omega_{\alpha}(y)\,\Delta_{\theta}(y)+\frac{\mu}{2}\mathbb{E}_{\pi_{\mathrm{ref}}}\big[\Delta_{\theta}(y)^{2}\big]}$ (8)
This retains the linear dynamics of the Euclidean geometry while remaining compatible with standard gradient-based training.
Remark 4.2 (On-Policy Anchoring).
We set $\pi_{\mathrm{ref}}=\pi_{\text{old}}$ (on-policy anchoring), ensuring the trust region condition $\Delta\approx 0$ holds locally, justifying the approximation.
4.6 Gradient Dynamics
For the ratio-coordinate objective Equation˜7, the gradient w.r.t. $v_{\theta}(y)$ is:
$\nabla_{v}\mathcal{L}_{\text{OPO}}^{\text{ratio}}=-\omega_{\alpha}(y)+\mu\,v_{\theta}(y)$ (9)
with the unique equilibrium:
$v^{*}(y)=\frac{\omega_{\alpha}(y)}{\mu}$ (10)
This linear relationship has several desirable properties:
  1. 1.
    Stable Equilibrium: The objective is strictly convex in $v_{\theta}(y)$ for fixed $\omega_{\alpha}(y)$, yielding a unique stationary point in the ratio coordinate.
  2. 2.
    No Saturation: Unlike sigmoid-based losses, the gradient magnitude does not vanish as $v$ grows.
  3. 3.
    Orthogonal Control: Changing $\alpha$ (sampling) affects only $\omega_{\alpha}$; changing $\mu$ (optimization) affects only the regularization strength. The two do not interact.
@@ 5 Theoretical Analysis
We formlize the decoupling of design axes and analyze the distinct regularization properties of OPO.
5.1 Decoupled Sampling and Optimization Geometry
We analyze the optimization dynamics in the function space of ratio outputs $v$. This perspective captures the intrinsic objective geometry. We first formalize the connection to parameter space dynamics.
@@ 5.1 Decoupled Sampling and Optimization Geometry
This structural decoupling guarantees stable, predictable convergence dynamics in the function space.
@@
Gradient descent in $v$-space on the OPO objective follows a linear dynamical system:
@@ 5.3 Quantitative Gradient Analysis
This guarantees a linear driving force that does not vanish arbitrarily, preventing the saturation observed in KL-based methods.
@@ 5.4 Algorithm and Flowchart
Current Policy $\pi_{\theta}$ Sampling $(x,y)$ Evaluation $A(y)$ OPO Decomposition 1. Ratio $r=\pi/\pi_{old}$ 2. Reweight $\omega_{\alpha}\leftarrow r^{\alpha}A$ 3. Orth. Penalty $\mu(r-1)^{2}$ Update $\theta$ Next Iteration
Figure 1: Flowchart of the Orthogonalized Policy Optimization (OPO) framework. The core innovation lies in the “OPO Decomposition” block (green), which separates the advantage signal (reweighted by $\alpha$) from the trust-region constraint (quadratic penalty $\chi^{2}$), ensuring linear non-saturating gradients.
Algorithm 1: Orthogonalized Policy Optimization (OPO)  Input: Initial policy $\pi_{\theta}$, hyperparameters $\alpha,\mu,\eta$ for each iteration do 1. Set Reference: $\pi_{\mathrm{ref}}\leftarrow\pi_{\theta}$ (on-policy anchoring) 2. Collect Rollouts: Sample $(x,y)\sim\pi_{\theta}$, compute rewards $R(x,y)$ 3. Compute Advantages: $A(y)\leftarrow R(y)-\bar{R}$ (group-normalized) 4. Compute Weights: $\omega_{\alpha}(y)\leftarrow A(y)\cdot[\text{reweight by }\alpha]$ (sampling geometry) 5. Compute Ratio: $v_{\theta}(y)\leftarrow\pi_{\theta}(y)/\pi_{\mathrm{ref}}(y)-1$ (or use log-approx) 6. Compute Loss: $\mathcal{L}\leftarrow-\sum_{y}\omega_{\alpha}(y)v_{\theta}(y)+\frac{\mu}{2}\sum_{y}v_{\theta}(y)^{2}$ (optimization geometry) 7. Update Policy: $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}$ end for
Figure 2: Pseudocode for Orthogonalized Policy Optimization. The algorithm explicitly decouples sampling geometry (step 4) from optimization geometry (step 6).
@@ 6.1 Experimental Setup
  • OPO: The proposed method with $\alpha=0.6$, $\mu=1.0$, on-policy anchoring ($\pi_{\mathrm{ref}}=\pi_{\text{old}}$), and adaptive $\tau$ with range $[0.2,1.5]$.
@@ 6.2 Results
Table 1: Comparison of alignment algorithms on Qwen3-1.7B + MATH Level 3. Reward and gradient norm are averaged over the final 20 training steps.
@@
Figure 4 illustrates the gradient norm dynamics across training. OPO exhibits consistently higher gradient norms, validating the non-saturating property of the $\chi^{2}$ geometry. DAPO shows severely diminished gradients, explaining its early plateau.
@@
Figure 5 depicts the policy entropy throughout training. OPO maintains consistently higher entropy levels compared to baselines. We attribute this to the granularity of supervision: GRPO and DAPO utilize token-level updates, imposing dense supervision that forces the policy to collapse rapidly into specific phrasings. In contrast, OPO (and GSPO) employs sequence-level updates. This “sparse” supervision aligns the total trajectory probability without micromanaging individual tokens, thereby preserving policy diversity and exploration potential.

### v3

@@ Abstract
We propose a principled structural remedy: formulate alignment as an orthogonal mirror descent problem, in which sampling geometry enters as a linear driving force derived from an $\alpha$-divergence projection, while optimization geometry is determined independently by a Bregman divergence (mirror map). This perspective leads to a new alignment objective—Orthogonalized Policy Optimization (OPO)—obtained by choosing a Euclidean mirror map in likelihood-ratio space. The resulting objective admits a closed-form solution, linear (non-saturating) gradient dynamics, and a well-conditioned trust region, while remaining fully compatible with standard LLM training pipelines.
@@ 1 Introduction
  • We propose the Orthogonal Mirror Descent framework, which decouples these design axes via two independent geometries: an $\alpha$-divergence projection governs sampling, while a Bregman divergence (mirror map) governs optimization.
@@ 2 Related Work
3 Preliminaries: Two Coordinate Systems
Before presenting the OPO framework, we establish the two coordinate systems that underlie all ratio-based alignment objectives.
3.1 Ratio Space vs. Log-Ratio Space
@@ 4.1 The General Framework
4.2 Axis 1: Sampling Geometry via $\alpha$-Divergence Projection
The sampling weight $\omega_{\alpha}$ determines the effective weighting of training examples. We derive it from a principled $\alpha$-divergence interpolation, providing the sampling axis with the same geometric rigor as the optimization axis.
Definition 4.1 (Oracle Target Distribution).
Given a non-negative quality signal $Q(y)\geq 0$ and the behavior policy $\tilde{p}(y):=\pi_{\mathrm{old}}(y)$, define the oracle target distribution:
$P^{*}(y)\;\propto\;Q(y)$ (4)
In the standard RL setting with advantage function $A(y)$, the natural choice is $Q(y)=\exp(A(y))$, corresponding to the Boltzmann-optimal density. The importance sampling (IS) ratio is $w_{\mathrm{IS}}(y)=P^{*}(y)/\tilde{p}(y)\propto Q(y)/\tilde{p}(y)$.
Direct use of $w_{\mathrm{IS}}$ suffers from variance explosion in large action spaces. The parameter $\alpha\in[0,1]$ geometrically tempers this correction by interpolating along the $\alpha$-geodesic [11] connecting $\tilde{p}$ to $P^{*}$:
Proposition 4.2 ($\alpha$-Escort Weighting).
The $\alpha$-geodesic between $\tilde{p}$ and $P^{*}$ yields the escort distribution $\rho_{\alpha}(y)\propto\tilde{p}(y)^{\alpha}\,P^{*}(y)^{1-\alpha}$. The corresponding importance weight is:
$\boxed{\omega_{\alpha}(y)\;=\;\frac{\rho_{\alpha}(y)}{\tilde{p}(y)}\;\propto\;\left(\frac{P^{*}(y)}{\tilde{p}(y)}\right)^{1-\alpha}=\left(\frac{\exp(A(y))}{\tilde{p}(y)}\right)^{1-\alpha}}$ (5)
This parameterization has clear geometric semantics:
  • $\alpha\to 1$: $\omega_{\alpha}\to 1$ (uniform weighting; mode-covering, conservative).
  • $\alpha\to 0$: $\omega_{\alpha}\to P^{*}/\tilde{p}$ (full IS correction; peak-seeking, aggressive).
  • $0<\alpha<1$: Smooth interpolation governed by $\alpha$-geometry.
Remark 4.3 (On-Policy Implementation with Frozen Denominator).
In the on-policy setting where $\tilde{p}=\pi_{\mathrm{old}}$, the theoretical escort weight $\omega_{\alpha}=(\exp(A)/\tilde{p})^{1-\alpha}$ can be decomposed as:
$\omega_{\alpha}(y)\;=\;\exp\!\big((1-\alpha)\,A(y)\big)\;\cdot\;\operatorname{sg}\!\left(\frac{1}{\pi_{\mathrm{old}}(y)^{\,1-\alpha}}\right)$ (6)
where $\operatorname{sg}(\cdot)$ denotes the stop-gradient operator. This formulation is not an approximation—it computes the exact escort weight from Eq. (5). The stop-gradient on the $1/\pi_{\mathrm{old}}^{1-\alpha}$ factor serves two critical purposes: (i) it prevents the $1/p$ denominator from injecting unbounded gradients, preserving the constant Hessian $\nabla^{2}_{v}\mathcal{L}=\mu I$; and (ii) it is semantically natural, since $\pi_{\mathrm{old}}$ is already frozen as the reference policy in each iteration.
Remark 4.4 (Two-Geometry Structure).
With this derivation, the entire OPO framework admits a clean geometric decomposition:
Alignment == $\alpha$-geometry (sampling) $\times\;$ mirror geometry (optimization)
The sampling axis is an $\alpha$-divergence interpolation in the statistical manifold; the optimization axis is a Bregman divergence projection in ratio space. These two geometries are structurally orthogonal: $\alpha$ shapes the driving force without altering the Hessian, and $\mu$ shapes the curvature without altering the target.
@@ 4.3 Axis 2: Optimization Geometry: Euclidean Mirror Map
This choice transforms the alignment problem into a quadratic optimization in ratio space, with curvature $\mu I$ independent of the data.
4.4 The OPO Objective
Combining the $\alpha$-escort sampling weight (Axis 1) with the Euclidean mirror map (Axis 2) yields the OPO objective:
@@ 4.4 The OPO Objective
where $\mu=1/\eta$ is the stiffness parameter, with closed-form solution $v^{*}(y)=\omega_{\alpha}(y)/\mu$.
4.5 Gradient Dynamics: Why OPO Does Not Saturate
Differentiating Eq. (9) w.r.t. $v_{\theta}(y)$:
$\nabla_{v}\mathcal{L}_{\text{OPO}}=-\omega_{\alpha}(y)+\mu\,v_{\theta}(y)$ (10)
with unique equilibrium $v^{*}(y)=\omega_{\alpha}(y)/\mu$. This linear relationship has three desirable properties:
  1. 1.
    Stable Equilibrium: Strict convexity in $v_{\theta}$ yields a unique stationary point.
  2. 2.
    No Saturation: Unlike sigmoid-based losses, the gradient does not vanish as $v$ grows—it increases proportionally to the distance from equilibrium.
  3. 3.
    Orthogonal Control: Changing $\alpha$ (sampling) affects only $\omega_{\alpha}$; changing $\mu$ (optimization) affects only the curvature. The two do not interact.
4.6 Implementation: Log-Ratio Approximation
The ratio $v_{\theta}=\pi_{\theta}/\pi_{\mathrm{ref}}-1$ is not directly available from LLM logits. However, for the log-ratio $\Delta_{\theta}=\log\pi_{\theta}-\log\pi_{\mathrm{ref}}$ (which is directly computable), the first-order Taylor expansion $v=e^{\Delta}-1\approx\Delta$ holds within the trust region. This yields the implementable form:
$\boxed{\mathcal{L}_{\text{OPO}}^{\text{log}}\approx-\mathbb{E}_{y\sim\pi_{\mathrm{ref}}}\big[\omega_{\alpha}(y)\,\Delta_{\theta}(y)\big]+\frac{\mu}{2}\mathbb{E}_{\pi_{\mathrm{ref}}}\big[\Delta_{\theta}(y)^{2}\big]}$ (11)
This retains the linear dynamics of Euclidean geometry while operating entirely in log-space. The approximation is justified by on-policy anchoring: setting $\pi_{\mathrm{ref}}=\pi_{\text{old}}$ ensures $\Delta\approx 0$ at the start of each iteration.
4.7 Constructive Derivation: Regression in Ratio Space
We provide an alternative, constructive derivation that arrives at the same OPO loss from pure first principles—no mirror maps required.
Step 1: Define the target offset.
Given the escort weight $\omega_{\alpha}$ and regularization strength $\mu$, the optimal ratio offset is $v^{*}(y):=\omega_{\alpha}(y)/\mu$. This is the ideal deviation from the reference: the quality signal drives the direction; $\mu$ bounds the magnitude.
Step 2: Minimize the distance.
The natural metric in ratio space is mean squared error:
$\mathcal{J}(\theta)=\frac{\mu}{2}\big(v_{\theta}(y)-v^{*}(y)\big)^{2}$ (12)
Step 3: Complete the square.
Expanding and discarding the $\theta$-independent constant $\omega_{\alpha}^{2}/(2\mu)$:
$\mathcal{J}(\theta)=\underbrace{\frac{\mu}{2}\,v_{\theta}^{2}}_{\text{quadratic regularizer}}-\underbrace{\omega_{\alpha}\cdot v_{\theta}}_{\text{linear drive}}+\text{const}\;\;\Longrightarrow\;\;\mathcal{L}(\theta)=-\omega_{\alpha}\,v_{\theta}+\frac{\mu}{2}\,v_{\theta}^{2}$ (13)
which is exactly the OPO objective (Eq. 9). The completing-the-square identity provides a constructive proof: MSE regression in ratio space is gradient-equivalent to the linear-plus-quadratic OPO loss.
Remark 4.5 (Equivalence of Perspectives).
[…]
@@ 4.7 Constructive Derivation: Regression in Ratio Space
A physical analogy: in quantum mechanics, superpositions are linear in amplitude ($\psi$), while measurements are quadratic in amplitude. Similarly, OPO operates in the ratio space $v$ (“amplitude”), where preference signals superpose linearly, and regularization is quadratic. KL-based methods operate in log-probability (“energy”) space, where strong signals lead to exponential collapse. The amplitude picture explains why OPO sustains non-saturating gradients in high-confidence regimes.
@@ 5 Theoretical Analysis
The preceding section derived OPO from two complementary perspectives (mirror descent and ratio regression). We now prove that OPO’s key structural properties—orthogonal control and global contraction—hold rigorously.
5.1 Formal Decoupling of Geometry Axes
All analysis is conducted in the function space of ratio outputs $v\in L^{2}(\pi_{\mathrm{ref}})$, which captures the intrinsic objective geometry. We first bridge to parameter space.
@@ 5.1 Formal Decoupling of Geometry Axes
Remark 5.3 (Scope of Orthogonality).
The structural orthogonality in Theorem 5.2 holds exactly in the ratio function space $L^{2}(\pi_{\mathrm{ref}})$. In parameter space $\theta$, interaction may arise through policy parametrization, since $\nabla_{\theta}v=\nabla_{\theta}(\exp(\Delta)-1)$ depends on the training trajectory influenced by $\alpha$. However, this coupling is a property of the parametrization, not the intrinsic objective curvature: the Hessian $\nabla^{2}_{v}\mathcal{L}=\mu I$ remains structurally independent of $\alpha$ regardless of the parameter mapping.
A natural concern is whether the stop-gradient operator in $\omega_{\alpha}$ (Remark 4.3) introduces bias. The following lemma shows it does not, at the anchor point:
Lemma 5.4 (Exactness of Surrogate Gradient at Anchor Point).
Let $\mathcal{L}_{\mathrm{true}}(\theta)$ be the exact objective without stop-gradient on $\omega_{\alpha}$, and $\mathcal{L}_{\mathrm{OPO}}(\theta)$ the surrogate with $\operatorname{sg}$. Under on-policy anchoring $\pi_{\mathrm{ref}}=\pi_{\mathrm{old}}$, at the start of each iteration ($\theta=\theta_{\mathrm{old}}$):
$\left.\nabla_{\theta}\mathcal{L}_{\mathrm{OPO}}(\theta)\right|_{\theta=\theta_{\mathrm{old}}}=\left.\nabla_{\theta}\mathcal{L}_{\mathrm{true}}(\theta)\right|_{\theta=\theta_{\mathrm{old}}}$ (17)
Furthermore, the gradient error grows only as $\mathcal{O}(\|\pi_{\theta}-\pi_{\mathrm{old}}\|)$.
Proof.
By the product rule, the true function-space gradient contains a cross-term $(\nabla_{v}\omega_{\alpha})\,v$ that the surrogate omits. At the anchor point $\pi_{\theta}=\pi_{\mathrm{old}}$, the ratio offset is exactly $v=\pi_{\theta}/\pi_{\mathrm{old}}-1=0$, so this cross-term vanishes. For $\theta\neq\theta_{\mathrm{old}}$, the error is bounded by $\|(\nabla_{v}\omega_{\alpha})\,v\|=\mathcal{O}(\|v\|)=\mathcal{O}(\|\pi_{\theta}-\pi_{\mathrm{old}}\|)$, which is small within the trust region. ∎
This confirms that stop-gradient is not an engineering hack but a first-order exact approximation: the surrogate and true objectives agree to first order at every on-policy anchor. Under sufficiently small trust-region updates (i.e., $\|v\|$ bounded by the $\chi^{2}$ penalty), the cumulative bias across iterations remains second-order small, since each anchor reset re-zeroes the first-order error term.
@@
Under the assumption that the policy class can represent $v^{*}$ and that local linearization of the map $\theta\mapsto v_{\theta}$ holds near the optimum, gradient descent in $v$-space on the OPO objective follows a linear dynamical system:
@@ 5.3 Quantitative Gradient Analysis
This guarantees a linear driving force that does not vanish arbitrarily. The practical consequence is stark: in mathematical reasoning, models frequently assign near-1.0 probability to correct intermediate steps. KL-based methods penalize these high-confidence correct outputs by exponentially shrinking the learning signal, whereas OPO maintains a constant reward-driven pull via its linear dynamics—allowing the model to continue refining even its most confident predictions.
@@ 5.4 Algorithm and Flowchart
Current Policy $\pi_{\theta}$ Sampling $(x,y)$ Evaluation $A(y)$ OPO Decomposition 1. Ratio $v=\pi/\pi_{old}-1$ 2. $\alpha$-Weight $\omega_{\alpha}=(e^{A}/\tilde{p})^{1-\alpha}$ 3. Orth. Penalty $\mu\,v^{2}$ Update $\theta$ Next Iteration
Figure 1: Flowchart of the Orthogonalized Policy Optimization (OPO) framework. The “OPO Decomposition” block (green) embodies the two-geometry structure: the $\alpha$-divergence projection (step 2) governs sampling, while the $\chi^{2}$ penalty (step 3) governs optimization. These two axes are structurally orthogonal.
Algorithm 1 Orthogonalized Policy Optimization (OPO)
1:Initial policy $\pi_{\theta}$; parameters $\alpha\in[0,1],\;\mu>0,\;\eta>0$
2:for each iteration do
3:  Anchor: $\pi_{\mathrm{ref}}\leftarrow\pi_{\theta}$ $\triangleright$ On-policy anchoring
4:  Rollout: Sample $(x,y)\sim\pi_{\theta}$; compute rewards $R(x,y)$
5:  Advantage: $A(y)\leftarrow\mathrm{Normalize}(R(x,y))$ $\triangleright$ Group-normalized
6:  Escort Weight: $\omega_{\alpha}(y)\leftarrow\exp\!\big((1{-}\alpha)\,A(y)\big)\cdot\operatorname{sg}\!\big(\pi_{\mathrm{ref}}(y)^{-(1-\alpha)}\big)$ $\triangleright$ Eq. 6
7:  Log-Ratio: $\Delta_{\theta}(y)\leftarrow\log\pi_{\theta}(y)-\log\pi_{\mathrm{ref}}(y)$
8:  Loss: $\mathcal{L}\leftarrow\frac{1}{N}\sum_{y}\big[-\omega_{\alpha}(y)\,\Delta_{\theta}(y)+\frac{\mu}{2}\,\Delta_{\theta}(y)^{2}\big]$ $\triangleright$ $y$: complete response sequence
9:  Update: $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}$
10:end for
@@ 6.1 Experimental Setup
  • OPO: The proposed method with $\alpha=0.4$, $\mu=1.0$, on-policy anchoring ($\pi_{\mathrm{ref}}=\pi_{\text{old}}$), and adaptive $\tau$ with range $[0.2,1.5]$. Note: $1-\alpha=0.6$ controls the effective advantage amplification in Eq. (6).
@@ 6.2 Results
Table 1: Comparison of alignment algorithms on Qwen3-1.7B + MATH Level 3. Reward and gradient norm are averaged over the final 20 training steps. Higher gradient norm at convergence, combined with stable (non-exploding) training dynamics, indicates absence of gradient saturation rather than optimization instability.
@@
Figure 3 illustrates the gradient norm dynamics across training. OPO exhibits consistently higher but stable (non-exploding) gradient norms throughout training, validating the non-saturating property of the $\chi^{2}$ geometry. The stability of these norms rules out the alternative explanation of optimization instability or under-regularization. DAPO shows severely diminished gradients, explaining its early plateau.
@@
Figure 4 depicts the policy entropy throughout training. OPO maintains consistently higher entropy levels compared to baselines. We attribute this to the granularity of supervision: GRPO and DAPO utilize token-level updates, imposing dense supervision that forces the policy to collapse rapidly into specific phrasings. In contrast, OPO (and GSPO) employs sequence-level updates, where $y$ denotes an entire response trajectory and the loss is averaged over complete sequences rather than individual tokens. This “sparse” supervision aligns the total trajectory probability without micromanaging individual tokens, thereby preserving policy diversity and exploration potential.